If the nominal interest rate per year is 16,5% per annum, compounded at the end of every second month, then the effective interest rate is

1.     18,00%.

2.     16,18%.

3.     14,53%.

4.     16,68%.

5.     17,68%.


1
Expert's answer
2021-05-21T12:48:34-0400

"r=(1+i/n)^n-1=(1+0.165/6)^6-1=0.1768=17.68\ \%"


Answer: 5. "17.68\ \%"